A drone with a mass 5.8 kg, is used to lift a mass , m= 3.2 kg by using an inextensible cable of negligible mass. The tension in the cable is T , and the drone accelerates vertically upwards at 0.9 m/s² . The lifting force of the drone is 180 N . The frictional force f experienced by the drone and the mass m , as they move upwards, is the same . Calculate the magnitude of the frictional force experienced by the drone and m
1
Expert's answer
2020-10-05T10:56:35-0400

F is the lifting force

f is the frictional force

M is the mass of the drone

m is the mass of the load

a is the acceleration.

Apply Newton's second law and assume that Y-axis looks upward:


"F-mg-Mg-f=(M+m)a,\\
f=F-(m+M)g-(M+m)a=\\
=F-(M+m)(g+a)=\\180-(5.8+3.2)(9.8+0.9)=83.7\text{ N}."